# Opinion Clustering under the Friedkin-Johnsen Model: Agreement in Disagreement

### v1

@@ Abstract
The convergence of opinions in the Friedkin-Johnsen (FJ) framework is well studied, but the topological conditions leading to opinion clustering remain less explored. To bridge this gap, we examine the role of topology in the emergence of opinion clusters within the network. The key contribution of the paper lies in the introduction of the notion of topologically prominent agents, referred to as Locally Topologically Persuasive (LTP) agents. Interestingly, each LTP agent is associated with a unique set of (non-influential) agents in its vicinity. Using them, we present conditions to obtain opinion clusters in the FJ framework in any arbitrarily connected digraph. A key advantage of the proposed result is that the resulting opinion clusters are independent of the edge weights and the stubbornness of the agents. Finally, we demonstrate using simulation results that, by suitably placing LTP agents, one can design networks that achieve any desired opinion clustering.
@@ 1 INTRODUCTION
In recent decades, social media networks have become a major platform for advertising, conducting socio-political campaigns, and even spreading misinformation. This evolution has generated a keen interest among researchers to study the impact of social interactions on the formation of public opinion. Several mathematical models have been proposed, such as the DeGroot’s model [1], the FJ model [2], the Hegselman-Krause model [3], and the Biased assimilation model [4], that explain complex aspects of human interactions such as individual biases and homophily. Among these, the FJ model is popular because it captures a diverse range of emergent behaviours. Its popularity also stems from its analytical tractability and performance over large datasets.
@@
The FJ model was proposed in [2] to justify the disagreement among closely interacting individuals. The opinions evolving under the FJ model converge to disagreement due to the presence of biased individuals, who are commonly known as the stubborn agents. In [2], the authors show that if each non-stubborn agent in the network has a directed path from a stubborn agent, the opinions converge asymptotically at a steady state. Further, they demonstrate that the final opinions depend on the initial opinions of only the stubborn agents and lie in the convex hull of these opinions. The authors in [5] introduced the notion of oblivious agents, who are non-stubborn and do not have a path from any stubborn agent. The algebraic and graph-theoretic conditions that ensure convergence in networks having oblivious agents were presented in [5] and [6], respectively. Additionally, in [6], the authors show that the final opinions depend on the initial opinions of both the stubborn agents and some oblivious agents.
In opinion dynamics, emergent behaviours such as consensus and opinion clustering are of keen interest to explain and achieve a desired social outcome. While consensus is being widely explored even in the FJ framework [7], this work focuses on opinion clustering. A network achieves opinion clustering when groups of two or more agents converge to the same opinion at a steady state. It represents swarm behaviour [8] and has applications in task distribution [9] and formation control [10]. The works [11, 12, 13] present graph-theoretic conditions under which opinion clustering occurs in continuous-time linear invariant systems. While it is known that disagreement occurs in the FJ framework, the conditions for opinion clustering are not well-explored. In [14], the authors present the graph-theoretic conditions to achieve opinion clustering in the FJ framework in networks with oblivious agents. However, the results apply only to those networks where each cycle contains only one stubborn agent.
Contrary to [14], our work aims to explain the emergence of opinion clusters in the FJ framework in arbitrarily connected digraphs. In opinion dynamics, an opinion cluster at steady state represents a group of individuals with a shared view on a topic. Our focus is specifically on how the network topology directs a certain group of agents to form an opinion cluster. With this, we now highlight our major contributions as follows:
@@
  • We introduce a new type of agent called the Locally Topologically Persuasive (LTP) agent. We show that each LTP agent is associated with a unique set of non-influential agents in its vicinity.
@@
  • Importantly, we show that a non-stubborn LTP agent can steer the opinions of those associated with it to form an opinion cluster. Thereby, highlighting that even non-stubborn agents can impact the final opinions.
The paper has been organised as follows: Sec.2 introduces the notations and the relevant preliminaries. Sec. 3 motivates the analysis of the topological dependence of opinion clusters. Sec. 4 presents a framework to relate final opinions of any set of agents in the graph. We introduce the LTP agents and present the conditions for opinion clustering in Sec. 5. Finally, we conclude in Sec. 6 with insights into future directions.
@@ 2.2 Graph Preliminaries
A walk is an ordered sequence of nodes such that each pair of consecutive nodes forms an edge in the graph. A cycle is a walk whose initial and final nodes coincide. If no nodes in a walk are repeated, it is a path. A graph is aperiodic if the lengths of all its cycles are coprime. An undirected graph is connected if a path exists between every pair of nodes. A directed graph (digraph) is a strongly connected graph if a directed path exists between every pair of nodes in the graph. A digraph is weakly connected if it is not strongly connected, but its undirected version is connected. A maximal subgraph of $\mathcal{G}$ which is strongly connected forms a strongly connected component (SCC). An SCC is an independent strongly connected component (iSCC) if every node in the SCC has all its in-neighbours within the same SCC.
2.3 Kron Reduction
Let $M\in\mathbb{R}^{p\times p}$ and $\alpha,\beta\subseteq[p]$ be index sets. Then $M[\alpha,\beta]$ is a submatrix of $M$ with rows indexed by $\alpha$ and columns indexed by $\beta$. The submatrix $M[\alpha]:=M[\alpha,\alpha]$ and $\alpha^{c}:=[p]\setminus\alpha$. If $M[\alpha^{c}]$ is non-singular, then the Schur complement of $M[\alpha^{c}]$ in $M$ is defined as:
[…]
Lemma 1 ([16])
  1. 1.
    $Q/\alpha^{c}$ is well defined if for each node $i\in\alpha^{c}$ there is a node $j\in\alpha$ such that a path $j\to i$ exists in $\mathcal{G}(Q)$. (Note that this condition is adapted to our framework because we consider $w_{ij}>0$ if edge $(j,i)\in\mathcal{E}$.)
  2. 2.
@@ 3 FJ model and emergent behaviours
Consider a network $\mathcal{G}$ of $n$ agents with vector $\mathbf{x}(k)=[x_{1}(k),...,x_{n}(k)]$ denoting the opinions of the agents at the $k^{th}$ instance. Under the FJ model, the opinions of agents evolve as follows:
@@
Suppose $\mathcal{G}$ is weakly connected and there is a non-stubborn agent who does not have a directed path from any stubborn agent. In [5], such agents are called oblivious agents. Without loss of generality, we can renumber the nodes in $\mathcal{G}$ such that the nodes $\{1,2,...,n_{o}\}$ are the oblivious agents and the rest are the non-oblivious, where $n_{o}\in[n]$. Additionally, we consider the oblivious agents that form an iSCC to be grouped together. Now, we can re-write the FJ-model (2) as:
@@
where $\mathbf{x}_{1}(k)\in\mathbb{R}^{n_{o}}$ and $\mathbf{x}_{2}(k)\in\mathbb{R}^{(n-n_{o})}$ denote the opinions of oblivious and non-oblivious agents. The matrix $W$ gets partitioned as $W=\begin{bmatrix}W_{11}&0\\ W_{21}&W_{22}\end{bmatrix}$ and $\bar{\beta}\in\mathbb{R}^{(n-n_{o})\times(n-n_{o})}$ is a diagonal matrix with entries equal to the stubbornness of the non-oblivious agents.
@@
Consider a weakly connected network $\mathcal{G}$ with opinions evolving under the FJ model (2).
  • If $\mathcal{G}$ does not contain any oblivious agents, the final opinions $\mathbf{x}^{*}$ always converge to,
    $\mathbf{x}^{*}=(I-(I-\beta)W)^{-1}\beta\mathbf{x}(0)$
  • If $\mathcal{G}$ has oblivious agents, then the opinions converge at steady state only if each iSCC composed of oblivious agents is aperiodic. The final opinions converge to,
@@
    where $W_{11}^{*}=\lim_{k\to\infty}W_{11}^{k}$.
Since $\mathbf{x}^{*}$ depends only on the initial opinions of the stubborn and the oblivious agents in iSCCs, these agents are called the influential agents. $\mathcal{I}$ denotes the set of all influential agents in $\mathcal{G}$. From Lemma 2, it follows that the impact of these influential agents on $\mathbf{x}^{*}$ depends on two factors: the network topology (captured by $W$) and the stubbornness of agents.
In general, the final opinions converge to different values, causing disagreement among agents. However, certain subgroups of agents can still have consensus amongst themselves and form opinion clusters in the final opinion. Naturally, the question arises how the formation of these opinion clusters depends on network topology and stubborn behaviour. The following example illustrates the effect of network topology on the opinion clusters.
@@
Consider the network $\mathcal{G}$ shown in Fig. 1(a). $\mathcal{G}$ has two stubborn agents: $2$ and $6$ with $\beta_{2}=0.3$, and $\beta_{6}=0.6$, and no oblivious agents. For initial opinions chosen from a uniform distribution on $[0,10]$, the opinions under the FJ model (2) evolve as shown in Fig. 1(b). The following opinion clusters form in the final opinion: agents $3,4$ and $5$ converge to $7.95$, agents $1,6$ converge to $8.85$ and agent $2$ converges to $6.83$. Next, we modify $\mathcal{G}$ by adding edge $(1,5)$ with weight $0.5$ (and reducing the weight of $(4,5)$ to $0.5$ to ensure $W$ remains row-stochastic) as shown in Fig.1(c). For the same stubbornness and the initial opinions, Fig. 1(d) shows that node $5$ forms a new cluster while the rest belong to the same opinion clusters.
@@
(c) Modified network $\mathcal{\hat{G}}$
(d) Evolution of opinions in $\mathcal{\hat{G}}$
Figure 1: Simple modification to $\mathcal{G}$ affects the opinion clusters in final opinion. Throughout the paper, we represent the stubborn agents and the non-stubborn agents by red nodes and orange nodes, respectively. The modified edges in $\mathcal{\hat{G}}$ are highlighted in red.
Example 1 demonstrates that the underlying topology has a significant impact on formation of opinion clusters in the network. Motivated by this, our objective is to establish the topology-based conditions that steer a subgroup of agents in the network to form an opinion cluster in the FJ framework.
4 The Relation of the final opinions with Network topology
In this section, we present a framework to relate the final opinions of agents with the underlying network topology. Consider a digraph $\mathcal{G}$ consisting of $n$ agents with opinions evolving under the FJ model (2). Let the network have $m$ stubborn agents. The final opinions satisfy the equations: $\mathbf{x}^{*}=(I-\beta)W\mathbf{x}^{*}+\beta\mathbf{x}(0)$. We can rewrite the above-mentioned linear equations as follows:
where $\mathbf{x}_{s}=[x_{s_{1}}(0),...,x_{s_{m}}(0)]$ and $s_{1},s_{2},...,s_{m}\in[m]$ are the labels of the stubborn agents. The matrix $\eta=[\eta_{ij}]\in\mathbb{R}^{n\times m}$ has entry $\eta_{{i}j}=\beta_{i}$ if the stubborn agent $i$ is labelled as $s_{j}$, otherwise $\eta_{ij}=0$ for $i\in[n],j\in[m]$. Henceforth, we represent eqn. (3) as $Rz=\mathbb{0}$ where $R=[r_{i,j}]\in\mathbb{R}^{(n+m)\times(n+m)}$.
@@ 4 The Relation of the final opinions with Network topology
The digraph $\mathcal{G}(R)$ can be induced from $R$ if an edge $(i,j)$ exists in $\mathcal{G}(R)$ when $r_{ji}<0$. The associated graph $\mathcal{G}(R)$ has $n+m$ nodes. Being consistent with the indexing of nodes in $\mathcal{G}$, the first $n$ nodes in $\mathcal{G}(R)$ are associated with $x_{i}^{*}$ for $i\in[n]$ and the nodes $n+1$ to $n+m$ are associated with the initial states of the $m$ stubborn agents. Note that each node $n+i$ forms a source in $\mathcal{G}(R)$ for $i\in[m]$. The source $n+i$ has a single outgoing edge to the node associated with the final opinion of the corresponding stubborn agent. For example, in Fig. 2, the initial opinion of stubborn agent $2$ is associated with node $n+1$ in $\mathcal{G}(R)$, thus, $n+1$ only has the outgoing edge $(n+1,2)$.
@@
Consider the network $\mathcal{G}$ shown in Fig. 1(a) with stubborn agents $2$ and $6$. The network $\mathcal{G}(R)$ for the matrix $R$ derived from eqn. (3) is shown in Fig. 2. The nodes in $\mathcal{G}(R)$ that represent the initial opinion of the stubborn agents are highlighted in turquoise.
Figure 2: The network $\mathcal{G}(R)$ derived from $\mathcal{G}$.
Remark 2
Since $R$ is a Laplacian matrix, it can be reduced by Kron reduction for a suitable choice of the subset of nodes $\alpha$ (defined in Sec. 2.3). Consider a network $\mathcal{G}$ without any oblivious agents. For such a network, each node in $[n]$ has a path in $\mathcal{G}(R)$ from at least one source $\{n+1,...,n+m\}$. Thus, by Statement 1) in Lemma 1, the Schur complement $R/\alpha^{c}$ is well-defined for any $\alpha$ such that $\{n+1,...,n+m\}\subseteq\alpha\subseteq[n+m]$. If $\mathcal{G}$ has an oblivious agent, then the sources $\{n+1,...,n+m\}$ do not have a path to this agent. Thus, we define $\alpha$ to include the oblivious agents belonging to each iSCC along with sources $\{n+1,...,n+m\}$. For this choice of $\alpha$, each $i\in\alpha^{c}$ has a path in $\mathcal{G}(R)$ from a node $j\in\alpha$. Thus, Statement 1) in Lemma 1 holds.
Importantly, eliminating the states associated with nodes in $\alpha^{c}$ in eqn. (3), reduces it to
where $\omega=\alpha\setminus\{n+1,...,n+m\}$. Thus, the Kron reduction technique is useful in determining the relations between the final opinions of a subgroup of nodes in $\mathcal{G}$ and quantifying their dependence on the influential agents (stubborn agents and oblivious agents in iSCC). Using the Kron-reduction technique, we present topology-based conditions under which the agents in a subgroup have equal final opinions and form an opinion cluster.
5 Topology-based opinion clustering
Consider a weakly connected network $\mathcal{G}$ with opinions evolving under the FJ model (2). Let $q$ be a non-influential agent in $\mathcal{G}$. Then, it is simple to see that the final opinion $x_{q}^{*}$ depends on the initial opinions of only those influential agents that have a path to $q$. Let $\mathcal{I}_{q}\subseteq\mathcal{I}$ be the set of all influential agents that have a directed path to $q$. Now, we define an LTP agent in terms of paths from agents in $\mathcal{I}_{q}$ to $q$.
Definition 1
Consider an agent $p$ in network $\mathcal{G}$. If every path in $\mathcal{G}$ from each influential agent $s\in\mathcal{I}_{q}$ to $q$ (where $p\neq q$) traverses $p$, then $p$ is an LTP agent and $q$ is said to be persuaded by $p$. Note that an LTP agent itself is not persuaded by any other agent, i.e. all paths from agents in $\mathcal{I}_{p}$ to $p$ must not pass through any other agent.
The set of all non-influential agents persuaded by an LTP agent $p$ is denoted as $\mathcal{N}_{p}$. Since $p$ is an LTP agent and $q\in\mathcal{N}_{p}$, it implies that all possible walks from each agent in $\mathcal{I}_{q}$ to $q$ also traverse $p$. An LTP agent can be an influential agent (i.e. a stubborn agent or an oblivious agent in an iSCC) or a non-influential agent. An influential agent is an LTP agent only if each path from the remaining influential agents in $\mathcal{I}_{q}$ to $q$ traverses $p$. The following example illustrates the identification of an LTP agent and the agent(s) it persuades.
Example 3
Consider the network $\mathcal{G}$ shown in Fig. 1(a). Since $\mathcal{G}$ is strongly connected, only the stubborn agents $2$ and $6$ are influential and have a path to all agents. Fig 3(a) demonstrates that each path from $2$ to $4$ (and $5$) in $\mathcal{G}$ traverses through $3$. Moreover, the paths from $6$ to $4$ (and $5$) also traverse through $3$ as displayed in Fig. 3(b). Since all paths from $2$ and $6$ to $3$ do not traverse any common node. Thus, $3$ is an LTP agent and $\mathcal{N}_{3}=\{4,5\}$. Similarly, $6$ is also an LTP agent and $\mathcal{N}_{6}=\{1\}$.
@@ 5 Topology-based opinion clustering
Figure 3: Node 33 is an LTP agent and it persuades 44 and $5$ as each path from the stubborn agents traverse $3$.
Suppose a network $\mathcal{G}$ has multiple LTP agents. The following question naturally arises: Can an agent be persuaded by two different LTP agents? The following Lemma presents an answer to this question.
@@
Under C1, every path from an agent in $\mathcal{I}_{q}$ to $j$ also traverses $i$. Since $j$ has a path to $q$, it follows that $\mathcal{I}_{j}\subseteq\mathcal{I}_{q}$, and hence every path from each node in $\mathcal{I}_{j}$ to $j$ traverses $i$. Thus $j\in\mathcal{N}_{i}$, so $j$ cannot be an LTP agent. Equivalently, under C2, we obtain $i\in\mathcal{N}_{j}$, so $i$ is not an LTP agent. Hence, both C1 and C2 result in contradiction.
@@
Under C3, there exists a path $P_{1}$ from $s$ to $q$ that traverses $i$ before $j$ and another path $P_{2}$ from $u$ to $q$ that traverses $j$ before $i$. We can represent $P_{1}$ as $s\to i\to j\to q$ and $P_{2}$ as $u\to j\to i\to q$. By definition, a path does not contain repeated nodes. It follows from $P_{1}$ that a path $j\to q$ exists that does not traverse $i$. Further, we know from $P_{2}$ that a path from $u\to j$ does not traverse $i$. Therefore, a path from $u\to q$ exists that does not traverse $i$. Consequently, $q\notin\mathcal{N}_{i}$ which contradicts our assumption that $q\in\mathcal{N}_{i}\cap\mathcal{N}_{j}$. Similarly, we can show that a path from $s\to q$ exists that does not traverse $j$. Since all of the conditions: C1, C2 and C3, lead to a contradiction, it means that the assumption $q\in\mathcal{N}_{i}\cap\mathcal{N}_{j}$ does not hold.

### v2

@@ Abstract
The convergence of opinions in the Friedkin-Johnsen (FJ) framework is well studied, but the topological conditions leading to opinion clustering remain less explored. To bridge this gap, we examine the role of topology in the emergence of opinion clusters within the network. The key contribution of the paper lies in the introduction of the notion of topologically prominent agents, referred to as Locally Topologically Persuasive (LTP) agents. Interestingly, each LTP agent is associated with a unique set of (non-influential) agents in its vicinity. Using them, we present conditions to obtain opinion clusters in the FJ framework in any arbitrarily connected digraph. A key advantage of the proposed result is that the resulting opinion clusters are independent of the edge weights and the stubbornness of the agents. Finally, we demonstrate that by suitably placing LTP agents, one can design networks that achieve any desired opinion clustering.
@@ 1 INTRODUCTION
The recent evolution of social media networks has generated a keen interest among researchers to study the impact of social interactions on public opinion. Several mathematical models such as the DeGroot’s model [1], the FJ model [2], the Hegselmann-Krause model [3], and the Biased assimilation model [4], explain complex aspects of human interactions such as individual biases and homophily. Among these, the FJ model is most analytically tractable and explains the frequently observed phenomenon of disagreement among closely interacting individuals [5]. Moreover, it is widely popular because it is empirically validated on real-world datasets [6], [7].
@@
The opinions evolving under the FJ model converge to disagreement due to the presence of biased individuals, who are called stubborn agents. In [2], the authors show that if each non-stubborn agent in the network is reachable from a stubborn agent, the final opinions depend on the initial opinions of only the stubborn agents. Reference [8] introduces the notion of oblivious agents, who are non-stubborn and do not have a path from any stubborn agent. The algebraic and graph-theoretic conditions that ensure convergence in networks with oblivious agents are presented in [8] and [9], respectively.
In opinion dynamics, emergent behaviours such as consensus, polarisation and opinion clustering are of interest to explain and achieve a desired social outcome. Several works [6, 10, 11, 12] have examined the FJ framework to reach these behaviours. In [6], the authors design stubbornness values that lead to behaviours such as consensus and polarisation. On the other hand, reference [10] presents topology-based conditions for graphs with two stubborn agents that result in polarisation. Furthermore, the authors in [11] examine the FJ model with randomised interactions and reveal that the opinions of non-stubborn agents can exhibit oscillations in presence of stubborn agents. In contrast, the present work focuses on opinion clustering, which is less-explored in the FJ framework. Opinion clustering occurs when a group of two or more agents converge to the same final opinion. It represents swarm behaviour and has applications in task distribution [13] and formation control [14]. In the sociological framework, opinion clustering represents the formation of groups with aligned views.
In this work, we examine the network topology towards achieving opinion clustering. Recent works [10, 11] in the FJ framework highlight the impact of topology in shaping emergent behaviours. More recently, the study of network topology is receiving attention to analyse the effects of recommender algorithms on opinion formation. The seminal works [15] and [16] contributed significantly by quantifying the influence of a stubborn agent using hitting probabilities of random walks over the network. These studies enabled the design of topology-based interventions (recommendations) for desirably shaping opinions [16, 17]. In the present work, we further explore the relation between the network topology and final opinions and address the following problem: Under what topology-based conditions, does a set of agents converge to the same final opinion (form an opinion cluster)? We leverage this relationship to obtain desired opinion clustering. In [12], the authors present the graph-theoretic conditions to achieve opinion clustering in the FJ framework in networks with oblivious agents. However, the result applies only to networks where each cycle contains at most one stubborn agent. Contrary to [12], our work aims to explain the emergence of opinion clusters in the FJ framework in arbitrarily connected digraphs. With this, we now highlight our major contributions:
@@
  • We define the notion of the Locally Topologically Persuasive (LTP) agent. Each LTP agent is associated with a unique set of non-influential agents in its vicinity.
@@
Paper Organisation: Sec.2 introduces the notations and the relevant preliminaries. Sec. 3 motivates the study of the formation of opinion clusters. Sec. 4, introduces the LTP agents and presents the conditions for opinion clustering. Finally, Sec. 5 concludes with insights into future directions.
@@ 2 Notations and Preliminaries
A walk is an ordered sequence of nodes such that each pair of consecutive nodes forms an edge in the graph. A cycle is a walk whose initial and final nodes coincide. A walk in which no node is repeated is a path. A path from node $x$ to node $y$ traverses node $z$ if the sequence of nodes that form this path contains $z$. A graph is aperiodic if the lengths of all its cycles are coprime. An undirected graph is connected if a path exists between every pair of nodes. A directed graph (digraph) is a strongly connected graph if a directed path exists between every pair of nodes in the graph. A digraph is weakly connected if the undirected version of the digraph is connected. A maximal subgraph of $\mathcal{G}$ which is strongly connected forms a strongly connected component (SCC). An SCC is an independent SCC (iSCC) if every node in the SCC has all its in-neighbours within the same SCC.
@@ 3 FJ model and emergent behaviours
Consider a network $\mathcal{G}$ of $n$ agents with vector $\mathbf{x}(k)=[x_{1}(k),...,x_{n}(k)]$ denoting the opinions of the agents at the $k^{th}$ instance. Under the FJ model, the opinions evolve as:
@@
Suppose $\mathcal{G}$ is weakly connected and there is a non-stubborn agent who does not have a directed path from any stubborn agent. Such agents are called oblivious agents [8]. Without loss of generality, we renumber the nodes in $\mathcal{G}$ such that the nodes $\{1,2,...,n_{o}\}$ are the oblivious agents and the rest are the non-oblivious, where $n_{o}\in[n]$. Additionally, we consider the oblivious agents that form an iSCC are grouped together. Now, we can re-write the FJ model (1) as:
@@
where $\mathbf{x}_{1}(k)\in\mathbb{R}^{n_{o}}$ and $\mathbf{x}_{2}(k)\in\mathbb{R}^{(n-n_{o})}$ denote the opinions of oblivious and non-oblivious agents. The matrix $W$ gets partitioned as $W=\begin{bmatrix}W_{11}&0\\ W_{21}&W_{22}\end{bmatrix}$ and the diagonal matrix $\bar{\beta}$ denotes the stubbornness of non-oblivious agents. Throughout this article, we consider that $\mathcal{G}$ satisfies the following:
Assumption 1
Each iSCC in $\mathcal{G}$ composed of the oblivious agents with two or more agents is aperiodic.
Assumption 1 is the necessary and sufficient condition for convergence under the FJ model [9].
@@
Consider a network $\mathcal{G}$ with $m\geq 1$ stubborn agents such that Assumption 1 holds. The opinions evolving under the FJ model (1) converge to,
@@
where $\mathbf{x}^{*}=[\mathbf{x}_{1}^{*},\mathbf{x}_{2}^{*}]$ denotes the final opinions and $W_{11}^{*}=\lim_{k\to\infty}W_{11}^{k}$.
From Lemma 1, it clearly follows that the final opinions depend on the initial opinions of only the stubborn agents and the oblivious agents in iSCCs. Hence, these agents are referred to as the influential agents. The set of all influential agents in $\mathcal{G}$ is denoted by $\mathcal{I}$.
The reference [9] establishes that the final opinions lie in the convex hull of the initial opinions of influential agents. In [15], the coefficients of each of these convex combinations (that determine the final opinions) are related to the network topology (captured by $W$) and the stubbornness of agents. Depending on these factors, the final opinions of agents (except the oblivious agents in an iSCC) generally converge to distinct values, which is termed as disagreement.
Disagreement is a more prevalent in social networks as compared to global consensus [6]. A special case of disagreement occurs when a subgroup of two or more agents converges to the same final opinion and forms an opinion cluster. An opinion cluster is defined as:
Definition 1
(Opinion Cluster) A set of agents $\mathcal{C}\subset\mathcal{V}$ form an opinion cluster if for any initial conditions $\mathbf{x}(0)\in\mathbb{R}^{n}$
$x_{i}^{*}-x_{j}^{*}=0\quad\forall\ i,j\in\mathcal{C}$
where $x_{k}^{*}$ denotes the final opinion of agent $k\in\mathcal{V}$.
From a sociological perspective, the existence of opinion clusters captures a more realistic distribution of opinions in society, where local consensus occurs in groups (echo chambers) despite widespread disagreement. The formation of opinion clusters has been examined in opinion dynamics [3], however, the role of network topology leading in this phenomenon remains less explored. In [12], the authors present topology-based conditions for the formation of opinion clusters in digraphs where each cycle has at most one stubborn agent. However, these conditions are insufficient to explain the emergence of opinion clusters in the digraph in Example 1 that contains a cycle with multiple stubborn agents.
@@
Consider a strongly connected network $\mathcal{G}$ in Fig. 1(a) with stubborn agents $2$ and $6$. The cycle in $\mathcal{G}$ formed by agents $1,2$ and $6$ contains two stubborn agents. Let the stubbornness of the stubborn agents be $\beta_{2}=0.3$, and $\beta_{6}=0.6$. For initial opinions from a uniform distribution on $[0,5]$, the final opinions under the FJ model (1) form three opinion clusters as shown in Fig. 1(b).
@@
Figure 1: Formation of opinion clusters. Throughout this paper, we represent the stubborn and non-stubborn agents by red and orange nodes, respectively.
In this work, our objective is to establish conditions under which a subgroup of agents in any arbitrarily connected digraph forms an opinion cluster in the FJ framework. Using these conditions, we can use network interactions as a tool to form desired opinion clusters.
@@
The agents within an opinion cluster converge to the same final opinions, regardless of initial opinions. This implies that each influential agent $s\in\mathcal{I}$ exerts an equal impact on all agents in the cluster. Thus, this work also presents the network structures under which a group of agents are equally affected by the influential agents.
4 Topology-based opinion clustering
Consider a weakly connected network $\mathcal{G}$ with opinions evolving under the FJ model (1). Let $q\in\mathcal{V}\setminus\mathcal{I}$ be a non-influential agent in $\mathcal{G}$. The final opinion $x_{q}^{*}$ depends on the initial opinions of only those influential agents that have a path to $q$. Let $\mathcal{I}_{q}\subseteq\mathcal{I}$ denote the set of influential agents that have a path to $q$. Certain topological properties in a network endow some agents with the special property of being ‘locally’ influential; we refer to them as LTP agents. More formally:
Definition 2
An agent $p\in\mathcal{V}$ is called an LTP agent if the following hold:
  1. (i)
    there exists an agent $q\in\mathcal{V}\setminus\mathcal{I}$ such that every path in $\mathcal{G}$ from each influential agent $s\in\mathcal{I}_{q}$ to $q$ (where $p\neq q$) traverses $p$,
  2. (ii)
    additionally, if $p\in\mathcal{V}\setminus\mathcal{I}$, then there does not exist an agent $c\in\mathcal{V}$ such that all paths from each agent $s\in\mathcal{I}_{p}$ to $p$ traverse $c$.
The non-influential agent $q$ is then said to be persuaded by $p$.
The set of all non-influential agents persuaded by an LTP agent $p$ is denoted by $\mathcal{N}_{p}$. If $p\in\mathcal{I}$ is an influential agent, condition (i) is sufficient to ensure that $p$ is an LTP agent. The condition (ii) in Defn. 2 is applicable only if $p$ is a non-influential agent. It conveys that a non-influential agent $p$, persuaded by an LTP agent $c$ (i.e $p\in\mathcal{N}_{c}$), is not an LTP agent itself. The following example illustrates which agents are LTP agents and which are the agents it persuades.
@@ 4 Topology-based opinion clustering
Consider the network $\mathcal{G}$ shown in Fig. 1(a). Since $\mathcal{G}$ is strongly connected, only the stubborn agents $2$ and $6$ are influential. Fig 2(a) demonstrates that each path from $2$ to $4$ (and $5$) in $\mathcal{G}$ traverses $3$. Further, the paths from $6$ to $4$ (and $5$) also traverse $3$ as shown in Fig. 2(b). Since all paths from $2$ and $6$ to $3$ do not traverse any common node. Thus, $3$ is an LTP agent and $\mathcal{N}_{3}=\{4,5\}$. Similarly, $6$ is also an LTP agent and $\mathcal{N}_{6}=\{1\}$.
Note that each path from $2$ and $6$ to $5$ traverses $4$. Hence, $4$ satisfies condition 1) in Definition 2 However, since $4\in\mathcal{N}_{3}$, it is not an LTP agent.
@@
Figure 2: Each path from the stubborn agents to 44 and 55 traverses $3$.
Remark 2
The notion of an LTP agent is analogous to that of cut vertices in undirected graphs (and strong articulation points in digraphs). A node $v$ is a cut vertex (or a strong articulation point in a strongly connected digraph) if and only if there are distinct nodes $x$ and $y$ such that every path from $x$ to $y$ traverses $v$ [19]. Since LTP agents are defined for digraphs, each LTP agent in a strongly connected digraph is a strong articulation point; however, the converse does not hold (e.g., node $4$ in Example 2).
The previous discussion describes the properties of the LTP agents. In larger networks, however, multiple LTP agents may exist. This leads to the question: Can a non-influential agent be persuaded by more than one LTP agent?
@@
Under C1, every path from an agent in $\mathcal{I}_{q}$ to $j$ also traverses $i$. Note that $j$ must be a non-influential agent under C1; otherwise $i$ will not be an LTP agent (as a path $j\to q$ exists that does not traverse $i$). Since $j$ has a path to $q$, it follows that $\mathcal{I}_{j}\subseteq\mathcal{I}_{q}$, and hence every path from each node in $\mathcal{I}_{j}$ to $j$ traverses $i$. Thus, $j\in\mathcal{N}_{i}$, so $j$ cannot be an LTP agent. Equivalently, under C2, we obtain $i\in\mathcal{N}_{j}$, so $i$ is not an LTP agent. Hence, both C1 and C2 result in contradiction.
@@
Under C3, there exists a path $P_{1}$ from $s$ to $q$ that traverses $i$ before $j$ and another path $P_{2}$ from $u$ to $q$ that traverses $j$ before $i$. We can represent $P_{1}$ as $s\to i\to j\to q$ and $P_{2}$ as $u\to j\to i\to q$. By definition, a path does not contain repeated nodes. It follows from $P_{1}$ that a path $j\to q$ exists that does not traverse $i$. Further, we know from $P_{2}$ that a path from $u\to j$ does not traverse $i$. Therefore, a path from $u\to q$ exists that does not traverse $i$. Consequently, $q\notin\mathcal{N}_{i}$ which contradicts our assumption that $q\in\mathcal{N}_{i}\cap\mathcal{N}_{j}$. Similarly, we can show that a path from $s\to q$ exists that does not traverse $j$. Since all of the conditions: C1, C2 and C3, lead to a contradiction, the assumption $q\in\mathcal{N}_{i}\cap\mathcal{N}_{j}$ does not hold.